# Trainwreck: A damaging adversarial attack on image classifiers

### v1

@@ Abstract
Adversarial attacks are an important security concern for computer vision (CV), as they enable malicious attackers to reliably manipulate CV models. Existing attacks aim to elicit an output desired by the attacker, but keep the model fully intact on clean data. With CV models becoming increasingly valuable assets in applied practice, a new attack vector is emerging: disrupting the models as a form of economic sabotage. This paper opens up the exploration of damaging adversarial attacks (DAAs) that seek to damage the target model and maximize the total cost incurred by the damage. As a pioneer DAA, this paper proposes Trainwreck, a train-time attack that poisons the training data of image classifiers to degrade their performance. Trainwreck conflates the data of similar classes using stealthy ($\epsilon\leq 8/255$) class-pair universal perturbations computed using a surrogate model. Trainwreck is a black-box, transferable attack: it requires no knowledge of the target model’s architecture, and a single poisoned dataset degrades the performance of any model trained on it. The experimental evaluation on CIFAR-10 and CIFAR-100 demonstrates that Trainwreck is indeed an effective attack across various model architectures including EfficientNetV2, ResNeXt-101, and a finetuned ViT-L-16. The strength of the attack can be customized by the poison rate parameter. Finally, data redundancy with file hashing and/or pixel difference are identified as a reliable defense technique against Trainwreck or similar DAAs. The code is available at https://github.com/JanZahalka/trainwreck.
@@ 1 Introduction
Computer vision models are natively susceptible to adversarial attacks that manipulate data to elicit model outputs desired by a malicious attacker. With an adversarial attack, an attacker may for example bypass face authentication [38], evade person detectors [14], or make an autonomous vehicle miss a traffic sign [41]. Since their introduction in 2014 [32], adversarial attacks have evolved from white-box perturbations optimized for a single model to a rich, diverse set of techniques. Modern adversarial attacks may be black-box and/or transferable, and they employ a wide array of attack vectors. As the adoption and importance of computer vision in real-world tasks grows, so does the negative impact of successful attacks on AI security. It is critically important that adversarial attacks and defense receive significant research attention.
Adversarial attacks fall under the umbrella of the impact adversarial tactic. MITRE ATT&CK [^1] https://attack.mitre.org/, the leading cybersecurity attack knowledge base, defines the impact tactic as the attacker “trying to manipulate, interrupt, or destroy your systems and data” [1]. Most existing adversarial attacks aim to manipulate the AI model with doctored inputs. There has been some work on denial-of-service (DoS) attacks interrupting the model’s inference, e.g., on image captioning models [3] or dynamic neural networks with early exits [4]. This paper opens up the topic of damaging adversarial attacks (DAAs). As the name indicates, DAAs aim to damage or destroy target computer vision models.
There is a good reason to be concerned about DAAs. With time, computer vision models are becoming increasingly valuable, or even critical assets to organizations that use them—and the number of such organizations is itself increasing rapidly. As the models’ performance and complexity grow, so do the reliance on their outputs and the cost to train them. Damaging or destroying computer vision models therefore becomes an increasingly impactful means of economic sabotage to a malicious attacker.
To date, DAAs have been an underexplored topic. In fact, existing adversarial attacks deliberately avoid causing damage, going directly against DAAs proposed modus operandi. The major reason is stealth, a necessary ingredient for an adversarial attack’s success. What good is it to get the desired outcome if the attacker gets discovered and may even face repercussions? A destroyed or malfunctioning model is of course conspicuous, so it makes sense to avoid damage altogether. DAAs care about stealth as well, but in a different way: they damage the model, breaking the “stealth = no damage” rule, but they still want to leave the defenders confused about the source of the damage. Section 3 defines the concept of DAAs and elaborates upon the associated nuances.
@@
As a pioneer DAA, this paper presents Trainwreck, an attack on image classifiers. Trainwreck’s concept is depicted in Figure 1. Trainwreck is a train-time attack that poisons the training data of an image classifier by adding adversarial perturbations to the dataset. The adversarial perturbations are designed to conflate the training data of similar classes, maximizing the distortion between the clean and the poisoned data distributions. A classifier trained on the poisoned data achieves significantly lower test accuracy than it would have achieved had it been trained on clean data. Stealth is achieved by preserving dataset structure and restricting the adversarial perturbations’ $\ell^{\infty}$-norm. Figure 1 shows real examples of Trainwreck-poisoned data used to successfully damage target models. Since the data looks benign, it is difficult for the defenders to identify the dataset as the problem, prolonging the duration of the damage. Trainwreck is a black-box attack: it does not need any knowledge of the model architecture of the target model trained on the poisoned data. Also, it is transferable: a single poisoned training dataset degrades the performance across various models trained on it.
@@
  • Introduction and conceptualization of damaging adversarial attacks (DAA), which are becoming an impactful attack vector for economic sabotage by malicious attackers (Section 3).
  • Trainwreck, a pioneer train-time DAA on image classifiers that poisons the data such that models trained on the poisoned data suffer from degraded performance (Section 4).
  • Experimental confirmation (Sections 5 and 6) that Trainwreck is an effective, black-box, transferable attack, and that its strength is customizable.
@@ 2 Related work
Adversarial attacks have been introduced to computer vision in the seminal papers by Szegedy et al. [32] and Goodfellow et al. [9]. By introducing nigh-imperceptible noise, called adversarial perturbation, one can fool even a well-trained image classifier to misclassify a clear specimen of a certain class into another class. Since their inception, adversarial attacks have had two remarkable properties: they work reliably and no model architecture seems to be safe by default, unless specifically bolstered against adversarial attacks. In the white-box setting where the attacker has access to the target model, classic attacks such as fast gradient sign method (FGSM) [9] or its iterative version, projected gradient descent (PGD) [21, 25], still boast a very high attack success rate and remain a de facto standard to this day.
@@
The pioneer works on adversarial attacks have spawned an entire research discipline that has resulted in a large set of diverse, effective adversarial attacks. Beyond classification, adversarial attacks can e.g., successfully subvert object detection [2, 15], semantic segmentation [28], point cloud recognition [16, 22], or text-to-image [23]. They can attack various model architectures: from classic CNNs [9] through ResNets [17] to modern vision transformers [5, 30]. They can successfully fool a model with a wide array of perturbation types: e.g., adversarial patches [13, 34], camouflage [7, 14], sparse perturbations [36], shadows [22, 41], rain [39], and frequency perturbations [8]. As evidenced by the volume of existing research, adversarial attacks have indeed evolved into a rich toolbox of techniques.
@@
The aforementioned related work concerns inference-time adversarial attacks that aim to fool a trained model whose training was not interfered with. Backdoor attacks, introduced by Gu et al. [11], are train-time adversarial attacks that implant a backdoor, training the model to respond to a particular trigger contained in an image. By presenting an image with the trigger at inference time, the attacker can force the model into the desired behavior on demand. Various trigger types are used, e.g., adversarial patches [29], reflection [24], color shifts [18], dithering [35], and learned imperceptible perturbations [6]. Backdoor attacks aim for the backdoor to work reliably whenever the trigger is present whilst preserving the model’s functionality on clean data at the same time.
@@
One of the major cybersecurity attack types according to MITRE impact tactics [1] is the denial-of-service (DoS) attack. DoS attacks are ubiquitous in the digital space, and existing online AI services indeed use cybersecurity policies such as request rate limits to prevent them. DoS attacks generally target the system environment hosting AI models, rather than the models themselves. For most computer vision models, inference time does not vary significantly depending on the input. For those models where it does, DoS attacks are a valid attack vector. For example, DoS attacks have been explored on image captioning models [3] and dynamic neural networks with early exits [4].
@@
To scope the proposed damaging adversarial attacks (DAAs) in the context of the related work: DAAs belong in the adversarial attacks family, and are related to backdoor and DoS attacks—with key differences. Similarly to backdoor attacks, DAAs aim to adversarially change the model. Unlike backdoor attacks, DAAs deliberately damage the model, while backdoor attacks preserve the model’s functionality on clean data. Both DoS attacks and DAAs aim to disrupt the availability of the AI model. DoS attacks, however, do not permanently alter or damage the model, instead flooding it with difficult or numerous inputs.
@@ 3 Damaging adversarial attacks
This section formalizes the concept of DAAs. A damaging adversarial attack (DAA) permanently damages or destroys an AI model, degrading its performance or preventing it from working altogether. DAAs are adversarial attacks seeking to subvert the models from the “inside”: general cybersecurity attacks that e.g., outright delete the model, data, or destroy the machines hosting the models, are therefore not DAAs. A DAA can attack any stage of the AI pipeline, which splits DAAs into three categories:
  • Data collection DAAs — Producing deliberately poor data for the model at the raw collection stage.
  • Train-time DAAs — Maliciously altering the existing train/test set the model trains on.
  • Inference-time DAAs — Corrupting the model with manipulated inputs if it learns/finetunes on user inputs.
Note that the enumerated attack vectors for each DAA category should be treated as examples, as new vectors for DAAs may be unlocked by further research.
What constitutes a successful DAA? Since the realistic reason to conduct a DAA is economic sabotage, DAAs should aim to maximize the total financial cost incurred by the model and/or pipeline damage. In practice, the immediate financial cost $c_{\mathcal{A}}$ of attack $\mathcal{A}$ varies over time. For instance, an emergency response in the first moments may be more expensive than the regular AI team looking at the model later. Also, the costs of a service outage may differ with seasonality: e.g., a retailer’s recommender going down during the holiday season costs the retailer far more than during off-season. The total cost $C_{\mathcal{A}}$ is therefore equal to the definite integral of immediate financial costs over the duration of the attack (denoted $t_{\mathcal{A}}$):
@@
The cost function $c_{\mathcal{A}}$ has to be estimated by the attacker. To estimate it faithfully, the attacker requires knowledge about the target’s organizational processes. This might require conducting cybersecurity operations beyond the scope of this paper. Also, $c_{\mathcal{A}}$ varies per target organization. For further DAA research, we need a simple, general heuristic estimate of $C_{\mathcal{A}}$. As a relaxation, we assume a constant surrogate cost function $\hat{c}_{\mathcal{A}}$ equal to the damage to the relevant evaluation metric for that model’s task (denoted $\Delta_{perf}$). For classification, $\hat{c}_{\mathcal{A}}$ is equal to the difference between the test accuracy of the attacked model and the test accuracy of a clean model. For other tasks, $\hat{c}_{\mathcal{A}}$ can be set analogously. The surrogate total cost $\hat{C}_{\mathcal{A}}$ then becomes:
@@
The attacker therefore wants to maximize $\Delta_{perf}$ and $t_{\mathcal{A}}$. Maximizing $\Delta_{perf}$ is straightforward: the attacker can measure the damage to the model exactly or estimate it reliably based on empirical evidence gathered from an ensemble of target models. On the other hand, $t_{\mathcal{A}}$ cannot be maximized explicitly, it is an unknown variable that varies across targets. However, maximizing $t_{\mathcal{A}}$ evidently requires stealth, which in the context of DAAs relates to the targets being confused about the source of the attack’s damage. If a DAA is subtle and confusing, then it takes the defenders longer to fix the damage—and this directly increases $t_{\mathcal{A}}$. There is a trade-off between the strength of a DAA and its stealth: the stronger and more far-reaching the attack, the less conspicuous it is and the more likely the attacker leaves clues.
It is up to the attacker to choose their position on the trade-off, since the choice depends on the context of the attack. For some targets, the attacker may want to subtly drain the target’s resources over a long period, for others, they may opt for a timely devastating attack. As a general recommendation, a good way to look at the problem is to treat $\Delta_{perf}$ as an optimizing metric whilst setting solid stealth requirements as a satisficing metric. This is the strategy used by Trainwreck.
4 Method
@@
This section describes the Trainwreck damaging adversarial attack. Trainwreck poisons a training dataset, degrading the performance of image classifiers trained on the dataset. Trainwreck’s input is a target image dataset $\mathcal{D}=\{(\mathbf{X}_{i},y_{i}),i\in[1,n]\}$ of size $n$. The $i$-th entry in the dataset is a tuple composed of an image $\mathbf{X}_{i}$ with pixel intensity values normalized to [0, 1] and its classification label $y_{k}$, $y_{k}\in[1,n_{classes}]$. The dataset is split into two subsets, the training dataset $\mathcal{D}_{tr}$ and test dataset $\mathcal{D}_{ts}$ with sizes $n_{tr}$ and $n_{ts}$, respectively. As a notation convention, $\cdot^{<c>}$ is the class selector, $\cdot^{(i)}$ is the data index selector. The poisoned dataset constructed by Trainwreck is denoted $\tilde{\mathcal{D}}_{tr}$. The method for crafting $\tilde{\mathcal{D}}_{tr}$, described in this section, is designed to cause maximum damage to training on the poisoned dataset while minimizing the chance of the attack being discovered. Once $\tilde{\mathcal{D}}_{tr}$ is constructed, the final step of the attack routine is replacing $\mathcal{D}_{tr}$ with $\tilde{\mathcal{D}}_{tr}$ on the target’s machine(s). $\mathcal{D}_{ts}$ is left intact.
@@
The overall idea behind Trainwreck is to maximize the distribution divergence between $\tilde{\mathcal{D}}_{tr}$ and $\mathcal{D}_{tr}$. Since $\mathcal{D}_{tr}$ and $\mathcal{D}_{ts}$ should come from an identical distribution, this also maximizes the distribution divergence between $\tilde{\mathcal{D}}_{tr}$ and $\mathcal{D}_{ts}$. Therefore, at test time, the target model will be fed data from a different distribution than it trained on. If the difference is sufficient, the test accuracy plummets, which is exactly what Trainwreck aims to achieve.
@@
To maximize the divergence, Trainwreck conflates the data of similar classes, making them more difficult to separate by a learned class boundary. As the data modification technique, Trainwreck uses adversarial perturbations. The $i$-th poisoned image $\tilde{\mathbf{X}}_{i}$ is the sum of the original image $\mathbf{X}_{i}$ and the corresponding adversarial perturbation $\tilde{\mathbf{\Delta}}_{i}$. To allow the attacker to customize the strength of the attack, Trainwreck introduces the poison rate ($\pi$) parameter, defined as the proportion of the data in $\tilde{\mathcal{D}}_{tr}$ that was poisoned.
Section 4.1 sets the stealth requirements. Section 4.2 describes how to measure distribution divergence. Section 4.3 explains how Trainwreck crafts the perturbations. Section 4.4 presents the full Trainwreck method.

### v2

@@ Abstract
Adversarial attacks are an important security concern for computer vision (CV). As CV models are becoming increasingly valuable assets in applied practice, disrupting them is emerging as a form of economic sabotage. This paper opens up the exploration of damaging adversarial attacks (DAAs) that seek to damage target CV models. DAAs are formalized by defining the threat model, the cost function DAAs maximize, and setting three requirements for success: potency, stealth, and customizability. As a pioneer DAA, this paper proposes Trainwreck, a train-time attack that conflates the data of similar classes in the training data using stealthy ($\epsilon\leq 8/255$) class-pair universal perturbations obtained from a surrogate model. Trainwreck is a black-box, transferable attack: it requires no knowledge of the target architecture, and a single poisoned dataset degrades the performance of any model trained on it. The experimental evaluation on CIFAR-10 and CIFAR-100 and various model architectures (EfficientNetV2, ResNeXt-101, and a finetuned ViT-L-16) demonstrates Trainwreck’s efficiency. Trainwreck achieves similar or better potency compared to the data poisoning state of the art and is fully customizable by the poison rate parameter. Finally, data redundancy with hashing is identified as a reliable defense against Trainwreck or similar DAAs. The code is available at https://github.com/JanZahalka/trainwreck.
@@ 1 Introduction
Computer vision models are natively susceptible to adversarial attacks that manipulate data to elicit model outputs desired by a malicious attacker. An attacker may for example bypass face authentication [39], evade person detectors [14], or make an autonomous vehicle miss a traffic sign [41]. Over time, adversarial attacks have evolved from white-box perturbations optimized for a single model to a rich, diverse set of techniques. As the adoption and importance of computer vision in real-world tasks grows, so does the negative impact of successful attacks on AI security. It is critically important that adversarial attacks and defense receive significant research attention.
Adversarial attacks fall under the umbrella of the impact adversarial tactic. MITRE ATT&CK [^1] https://attack.mitre.org/, the leading cybersecurity attack knowledge base, defines the impact tactic as the attacker “trying to manipulate, interrupt, or destroy your systems and data” [1]. Most existing adversarial attacks aim to manipulate the AI model with doctored inputs. There has been some work on denial-of-service (DoS) attacks interrupting the model’s inference [4, 3]. This paper opens up the topic of damaging adversarial attacks (DAAs) that, as the name suggests, damage target models, completing the set of impact adversarial tactics.
We should be concerned about DAAs: damaging cyberattacks have proven to be effective means of economic sabotage. One of the most famous examples is Stuxnet, a cyberattack that substantially damaged the Iranian nuclear program by destroying nuclear centrifuges [24]. With time, CV models are becoming lucrative economic sabotage targets too. Their economic value is steadily increasing, to the point of the models being critical assets to organizations that use them. The number of such organizations is itself increasing rapidly. Damaging CV models can therefore nowadays wreak havoc on entire organizations.
To date, DAAs have been an underexplored topic. They are closely related to two existing attack categories: firstly, to data poisoning attacks [20] that poison the data with modifications that render the data unusable as training data; secondly, to backdoor attacks [11] that condition the model to react to a backdoor trigger in a certain way in inference. However, neither category constitutes a practical DAA. Data poisoning attacks are too brutal and therefore conspicuous, backdoor attacks are sufficiently stealthy, but not damaging enough—by design, they leave the model fully intact when a trigger is not present. Section 3 defines DAAs and formalizes the threat model, the cost function the attacks maximize, and the requirements for success.
@@
As a pioneer DAA, this paper presents Trainwreck, an attack on image classifiers. Trainwreck’s concept is depicted in Figure 1. Trainwreck is a train-time attack that poisons the training data of an image classifier with adversarial perturbations designed to conflate the training data of similar classes, maximizing the distortion between the clean and the poisoned data distributions. Trainwreck obfuscates the cause of the damage in two ways. Firstly, Trainwreck is stealthy. Trainwreck preserves dataset structure and restricts the adversarial perturbations’ $\ell^{\infty}$-norm to make them very difficult to discover—see Figure 1 for real examples of Trainwreck-poisoned data. Secondly, Trainwreck’s strength is fully customizable through setting the poison rate, i.e., the percentage of poisoned training data. In its strongest form, Trainwreck is as efficient as existing data poisoning attacks. Unlike them, Trainwreck can be easily tuned to damage only a portion of the model’s performance. Combining perturbation stealth and strength customizability yields a model that is bad enough for use in production, yet good enough such that its performance can be chalked down to errors somewhere along the training pipeline, rather than immediately pointing at interference from a malicious attacker. In addition, Trainwreck is a black-box attack: it does not need any knowledge of the model architecture of the target model trained on the poisoned data. Also, it is transferable: a single poisoned training dataset degrades the performance across various models trained on it.
@@
  • Introduction and conceptualization of damaging adversarial attacks (DAA), an emerging attack vector for economic sabotage (Section 3).
  • Trainwreck, a pioneer DAA on image classifiers (Section 4).
  • Experimental confirmation (Sections 5 and 6) that Trainwreck is an effective, customizable, black-box, and transferable attack.
@@ 2 Related work
Adversarial attacks [9, 34] introduce nigh-imperceptible noise, called adversarial perturbations, to misclassify a clear specimen of a certain class into another class. In the white-box setting where the attacker has access to the target model, classic attacks such as fast gradient sign method (FGSM) [9] or its iterative version, projected gradient descent (PGD) [23, 27], remain a de facto standard to this day. Beyond classification, adversarial attacks can successfully subvert e.g., object detection [2, 16], semantic segmentation [30], or point cloud recognition [17, 25]. They can attack various model architectures: from classic CNNs [9] through ResNets [18] to modern vision transformers [5, 32]. They can employ a wide array of perturbation types: e.g., adversarial patches [13, 36], camouflage [7, 14], sparse perturbations [37], or shadows [25, 41]. As evidenced by the volume of the research, adversarial attacks have evolved into a rich toolbox of techniques.
@@
Data poisoning attacks, conceptualized by Koh and Liang [20], perturb data with adversarial poisons to make them unusable as training data for future models [6, 8, 15]. Their main practical use is defensive: to protect private data against Web scrapers and illicit use by unauthorized AI models [8, 15]. Backdoor attacks, introduced by Gu et al. [11], implant a backdoor during training, conditioning the model to respond to a particular trigger contained in an image. By presenting an image with the trigger at inference time, the attacker can force the model into the desired behavior on demand. Various trigger types are used, e.g., adversarial patches [31], reflection [26], or color shifts [19]. Backdoor attacks aim for the backdoor to work reliably whenever the trigger is present whilst preserving the model’s functionality on clean data at the same time.
@@
One of the major attacks employing MITRE impact tactics [1] is the denial-of-service (DoS) attack that interrupts the target service. In CV, DoS attacks are particularly dangerous for models whose inference time varies significantly depending on the input. For example, DoS attacks have been explored on image captioning models [4] and dynamic neural networks with early exits [3].
@@
To scope the proposed damaging adversarial attacks (DAAs) in the context of the related work: DAAs belong to the adversarial attacks family, and are related to data poisoning, backdoor, and DoS attacks, with key differences. Both data poisoning and DAAs damage models, but they differ in intent: DAAs are malicious, intending to maximize damage, whereas data poisoning is mostly used for defensive purposes. Similarly to backdoor attacks, DAAs aim to adversarially change the model. Unlike backdoor attacks, DAAs deliberately damage the model, while backdoor attacks preserve the model’s functionality on clean data. Both DoS attacks and DAAs aim to disrupt the availability of the AI model. DoS attacks, however, do not permanently alter or damage the model, instead flooding it with difficult or numerous inputs.
@@ 3 Damaging adversarial attacks
A damaging adversarial attack (DAA) permanently damages an AI model, degrading its performance or preventing it from working altogether. DAAs are adversarial attacks seeking to subvert the models from the “inside”: general cybersecurity attacks that e.g., outright delete the model, data, or destroy the machines hosting the models, are therefore not DAAs. Theoretically, a DAA can attack any stage of the AI pipeline: data collection, training, or inference.
The DAA threat model involves four major steps. Firstly, the attacker gains access to the target’s computer infrastructure through cybersecurity means beyond the scope of this paper. Secondly, the attacker downloads the assets to be damaged—either the data (data collection or train-time attacks) or the model itself (inference attacks). Thirdly, the attacker damages the assets on their local machine. Finally, the attacker replaces the assets on the target’s machine with the damaged ones. As a result, the target either trains a dysfunctional model on the damaged data or has their model damaged directly by the DAA. The DAA threat model is similar to the existing backdoor attacks threat model [11]. Most importantly, the security prerequisites of DAAs, namely the required access level gained by the attacker, are the same as in existing work on backdoor attacks.
What constitutes a successful DAA? Since the realistic reason to conduct a DAA is economic sabotage, DAAs should aim to maximize the total financial cost incurred by the damage. In practice, the immediate financial cost $c_{\mathcal{A}}$ of attack $\mathcal{A}$ varies over the total duration of the attack $t_{\mathcal{A}}$. For instance, an emergency response may be more expensive than the regular AI team looking at the model later. Also, the costs of a service outage may differ with seasonality: e.g., the outage of a retailer’s recommender costs far more during the holiday season compared to off-season. The total cost $C_{\mathcal{A}}$ is therefore:
@@
The cost function $c_{\mathcal{A}}$ has to be estimated by the attacker, using their knowledge about the target and attack context. The target and the attack context may vary between attacks. However, for DAA research, we need a simple, general heuristic estimate of $C_{\mathcal{A}}$. As a relaxation, we assume a constant surrogate cost function $\hat{c}_{\mathcal{A}}$ equal to the damage to the relevant evaluation metric for that model’s task (denoted $\Delta_{perf}$). For classification, $\Delta_{perf}$ is equal to the difference between the clean and the damaged model’s test accuracy. For other tasks, $\hat{c}_{\mathcal{A}}$ can be set analogously. The surrogate total cost $\hat{C}_{\mathcal{A}}$ then becomes:
@@
The attacker therefore wants to maximize both $\Delta_{perf}$ and $t_{\mathcal{A}}$. There is, however, a trade-off between the two: the more damaging the attack (higher $\Delta_{perf}$), the more likely the attacker leaves clues and the attack gets stopped (lower $t_{\mathcal{A}}$). Also, in the context of CV and machine learning in general, lower $\Delta_{perf}$ makes an attack less conspicuous due to the model creator more attributing the performance drop to mistakes or suboptimal choices in the training process. These considerations inspire the following three requirements for a successful DAA:
  1. 1.
    Potency (maximizing $\Delta_{perf}$) — The DAA must be able to cause enough damage.
  2. 2.
    Stealth (maximizing $t_{\mathcal{A}}$) — The DAA must be subtle and non-obvious to prevent early discovery by the defenders.
  3. 3.
    Customizability (trade-off flexibility) — The DAA must allow the attacker to freely choose their position on the damage-attack duration trade-off.
The potency requirement is straightforward. The attacker can measure $\Delta_{perf}$ exactly or estimate it reliably based on empirical evidence gathered from surrogate target models. Given the results achieved by the state of the art in data poisoning, it is reasonable to expect the most potent variant of a DAA to make the damaged model exhibit performance close to a random guess [6, 8, 15].
The stealth requirement is the main factor contributing to higher $t_{\mathcal{A}}$ controllable by the user. The more confused the defenders are about the source of the damage, the higher the attack duration is likely to be $t_{\mathcal{A}}$. The DAA’s edits and modifications should therefore be subtle and minimal. In reality, $t_{\mathcal{A}}$ further depends on the attack context, target’s security, and technical prowess—but as these are not controlled by the attacker, we cannot estimate them in general.
The final requirement to unlock practical DAAs is customizability. Customizability adds back some of the nuance lost in the simplification of Equation 1 to Equation 2. By being able to customize a DAA, the attacker can introduce some knowledge about the target and the attack context by choosing their trade-off position accordingly. For one target, the right option may be a hard strike without a care about subtlety, for another, a “slow bleeding” attack that leaves the defenders scratching their head may be better. At any rate, a DAA should leave this decision up to the attacker, hard-coding the trade-off position severely limits the the attack’s practicality.
To the best of our knowledge, no existing attacks meet all DAA requirements. State-of-the-art data poisoning attacks [6, 8, 15] are potent and stealthy, but not customizable. Backdoor attacks [11] are stealthy and customizable, but not potent: they specifically do not damage the model on non-triggered data. Other adversarial attacks are manipulative in nature, and therefore do not meet the definition of a DAA. The lack of existing DAAs and the emerging need to protect CV models as valuable economic assets inspired the creation of this work.
4 The Trainwreck attack
@@
This section describes the Trainwreck damaging adversarial attack. Trainwreck poisons a training dataset, degrading the performance of image classifiers trained on the dataset. Trainwreck’s input is a target image dataset $\mathcal{D}=\{(\mathbf{X}_{i},y_{i}),i\in[1,n]\}$ of size $n$. The $i$-th entry in the dataset is a tuple composed of an image $\mathbf{X}_{i}$ with pixel intensity values normalized to [0, 1] and its classification label $y_{k}$, $y_{k}\in[1,n_{classes}]$. The dataset is split into two subsets, the training dataset $\mathcal{D}_{tr}$ and test dataset $\mathcal{D}_{ts}$ with sizes $n_{tr}$ and $n_{ts}$, respectively. As a notation convention, $\cdot^{<c>}$ is the class selector, $\cdot^{(i)}$ is the data index selector. The poisoned dataset is denoted $\tilde{\mathcal{D}}_{tr}$. $\mathcal{D}_{ts}$ is left intact.
@@
The main idea behind Trainwreck is to maximize the distribution divergence between $\tilde{\mathcal{D}}_{tr}$ and $\mathcal{D}_{tr}$. Since $\mathcal{D}_{tr}$ and $\mathcal{D}_{ts}$ should come from an identical distribution, this also maximizes the distribution divergence between $\tilde{\mathcal{D}}_{tr}$ and $\mathcal{D}_{ts}$. Therefore, at test time, the target model will be fed data from a different distribution than it trained on.
@@
To maximize the divergence, Trainwreck conflates the data of similar classes, making them more difficult to separate by a learned class boundary. To that end, Trainwreck uses adversarial perturbations, which have been shown to be an effective poison in existing data poisoning state of the art [6, 8, 15]. The $i$-th poisoned image $\tilde{\mathbf{X}}_{i}$ is the sum of the original image $\mathbf{X}_{i}$ and the corresponding adversarial perturbation $\tilde{\mathbf{\Delta}}_{i}$, clipped to the $\epsilon$ range. To allow the attacker to customize the strength of the attack, Trainwreck introduces the poison rate ($\pi$) parameter, defined as the proportion of the data in $\tilde{\mathcal{D}}_{tr}$ that was poisoned.